OpenStudy (anonymous):

verify: tanxcscxcosx = 1

OpenStudy (whpalmer4):

Do you remember the definition of \(\csc x\)? How about \(\tan x\)?

OpenStudy (anonymous):

yes. secx and cotx

OpenStudy (anonymous):

1/secx and 1/cotx

OpenStudy (whpalmer4):

isn't \(\csc x = \dfrac{1}{\sin x}\)?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Also, isn't \[\tan x = \dfrac{\sin x}{\cos x}\]? What if you rewrite your equation using those substitutions?

OpenStudy (anonymous):

then it becomes sinx/cosx * cscxcosx =1

OpenStudy (whpalmer4):

put in the definition of csc x as well

OpenStudy (anonymous):

1=1!!

OpenStudy (whpalmer4):

looks like we verified it, eh?